OpenStudy (anonymous):

Three person A, B, C went to steal mangoes of neighbor. They collected as much mangoes as they could and dug and hid all the mangoes. A thought that B and C could betray him so, he went to place where they had hidden mangoes. Firstly, he took half of the mangoes and thought it was unfair so kept one back to the hole. B also thought same, so he went back and took half of the remaining mangoes. In the morning, C went and took all the mangoes. But, interestingly, they all took same number of mangoes. How many mangoes they had stolen that night? Three person A, B, C went to steal mangoes of neighbor. They collected as much mangoes as they could and dug and hid all the mangoes. A thought that B and C could betray him so, he went to place where they had hidden mangoes. Firstly, he took half of the mangoes and thought it was unfair so kept one back to the hole. B also thought same, so he went back and took half of the remaining mangoes. In the morning, C went and took all the mangoes. But, interestingly, they all took same number of mangoes. How many mangoes they had stolen that night? @Mathematics

OpenStudy (trancenova):

My boyfriend worked it out! They had stolen 6 mangoes!

OpenStudy (anonymous):

Yes! that's the answer.............. so, your boyfriend is mathematics genius

OpenStudy (trancenova):

A took 3, but put one back so he had 2 and there were 4 in the hole. B and C took half, 2 each. Yes yes he is, he's a logic powerhouse :P

OpenStudy (anonymous):

6 mangos: 6 / 2 = 3 3 - 1 = 2 4 / 2 = 2

OpenStudy (anonymous):

yup - good thinking!